Find the distance between (-4,0) and (-5,-3)
It is multiple choice:
A) 3 sqrt 10
B) 2
C) 6 sqrt 2
D) sqrt 10
How do I find the distance between the points? Thank you for your help!

Answer by venugopalramana(3286)

SEE THE FOLLOWING AND DO
-------------------------------------
Find the distance between (-3, -2)..CALL IT P SAY WITH (X1,Y1) COORDINATES
SO X1=-3...Y1=-2
and
(1, 4)...CALL IT Q SAY WITH (X2,Y2) COORDINATES.SO X2=1...Y2=4
DISTANCE (D)BETWEEN 2 POINTS PQ IS GIVEN BY FORMULA
D=SQRT[(X2-X1)^2+(Y2-Y1)^2]=SQRT[(1+3)^2+(4+2)^2]=SQRT[4^2+6^2]=SQRT(16+36)
=SQRT(52)=SQRT(4*13)=[SQRT(4)]=2*SQRT(13)............D IS THE ANSWER
A) sqrt 10
B) 27
C) 2 sqrt 5
D) 2 sqrt 13
Can anyone help me understand this problem? Thank you so very much!!